7.61 What is the de Broglie wavelength of a 145-g baseball traveling at 30.0 m/s (67.1 mph)? Is the wavelength much smaller o please be clear with the steps to solve and thanks so much!
Answer #1

m = 145g

      = 0.145Kg

v   = 30m/s

h   = 6.625*10^-34 J-sec

\lambda   = h/mv

     = 6.625*10^-34/(0.145*30)

      = 1.52*10^-34 m

The wave length is much smaller than diameter of of an atom [100Pm = 10^-10m ]